OpenStudy (anonymous):

Simplify 1/4(g − 16)

OpenStudy (anonymous):

g/4 - 4

OpenStudy (anonymous):

just multiply out

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

so a rule of multiplication is that if you have something like this (1/4)(g-16) you can multiply each thing together seperate. SO a(b-c) = ab -ac

OpenStudy (anonymous):

So, 1/4 *g = 1/4g and 1/4 * 16 = 4?

OpenStudy (anonymous):

yep